graph each system one solution no solution many solution
4x-2y=-6
2x-y=-3

Answer by Fombitz(32388)

1.4x-2y=-6
2.2x-y=-3
First graph equation 1.
When x=0, -2y=-6, y=3.
When y=0, 4x=-6, x=-3/2

Next, equation 2.
2.2x-y=-3
When x=0, -y=-3, y=3.
When y=0, 2x=-3, x=-3/2.
As you see the x and y intercepts are the same for equations 1 and 2.
Equations 1 and 2 are the same equation.
If you multiply equation 2 by (2), you get equation 1.
They have infinitely many solutions.